# How informative are summaries of the cosmic 21-cm signal?

### v1

@@
The cosmic 21-cm signal will bring data-driven advances to studies of the Cosmic Dawn (CD) and Epoch of Reionization (EoR). Radio telescopes such as the Square Kilometre Array (SKA) will eventually map the HI fluctuations over the first billion years – the majority of our observable Universe. With such large data volumes, it becomes increasingly important to develop ?optimal? summary statistics, allowing us to learn as much as possible about the CD and EoR. In this work we compare the astrophysical parameter constraining power of several 21-cm summary statistics, using the determinant of the Fisher information matrix, $\det\bm{F}$. Since we do not have an established ?fiducial? model for the astrophysics of the first galaxies, we compute for each summary the distribution of $\det\bm{F}$ across the prior volume. Using a large database of cosmic 21-cm lightcones that include realizations of telescope noise, we compare the following summaries: (i) the spherically-averaged power spectrum (1DPS), (ii) the cylindrically-averaged power spectrum (2DPS), (iii) the 2D Wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; (v) an information-maximizing neural network (IMNN); and (vi) the combination of 2DPS and IMNN. Our best performing individual summary is the 2DPS, having relatively high Fisher information throughout parameter space. Although capable of achieving the highest Fisher information for some parameter choices, the IMNN does not generalize well, resulting in a broad distribution across the prior volume. Our best results are achieved with the concatenation of the 2DPS and IMNN. The combination of only these two complimentary summaries reduces the recovered parameter variances on average by factors of $\sim$6.5 – 9.5, compared with using each summary independently. Finally, we point out that that the common assumption of a constant covariance matrix when doing Fisher forecasts using 21-cm summaries can significantly underestimate parameter constraints.
How informative are summaries of the cosmic 21-cm signal?
@@ 1 Introduction
The cosmic 21-cm signal, corresponding to the spin flip transition of the ground state of HI, provides a window on the first billion years of the Universe’s evolution. This under-explored period witnessed fundamental cosmic milestones, including the Cosmic Dawn (CD) of the first galaxies and the final phase change of our Universe: the Epoch of Reionization (EoR). Current radio interferometers are putting increasingly tighter upper limits on its power spectrum (PS; e.g. Trott et al. 2020; Mertens et al. 2020; HERA Collaboration et al. 2023), while the upcoming Square Kilometre Array (SKA) [^1] https://www.skatelescope.org is promising to provide a 3D image of the EoR in the next decade(s) (e.g. Koopmans et al. 2015; Mesinger 2020). The unknown properties of the first galaxies and IGM structures are encoded in the timing and morphology of the cosmic 21-cm signal (e.g. Pritchard & Furlanetto 2007; McQuinn & O’Leary 2012; Visbal et al. 2012; Pacucci et al. 2014).
@@
How do we infer these properties? The only statistically-robust way is through Bayesian inference: sampling from the theory and comparing forward models to the observation(s). However what should we use to compare the forward model to the data; i.e. how should we construct the likelihood? The cosmic 21-cm signal is intrinsically a 3D lightcone map. Performing inference directly on the lightcone would be incredibly challenging, due to the high-dimensionality of the data. Effectively, at each frequency the SKA should obtain a sky map comparable to current CMB maps, but with the full lightcone including thousands of such frequency slices (e.g. Loeb & Zaldarriaga 2004). [^2] Inference on high-dimensional lightcones is still possible in some cases, for example galaxy large scale structure (LSS) maps (e.g. Kitaura & Enßlin 2008; Jasche & Kitaura 2010; Jasche & Wandelt 2012, 2013; Leclercq et al. 2017; McAlpine et al. 2022; Dai & Seljak 2022; Bayer et al. 2022). Although such studies still compress the full galaxy observations, only treating them as biased tracers of the matter field, they are able to recover the phase information of our Universe’s initial conditions (so-called ?constrained realizations?). It remains to be seen if such studies can be extended to the Epoch of Reionization and Cosmic Dawn (though see Zhao et al. 2023b for a recent proof-of-concept study). Instead, the data is compressed into a summary statistic. Summary statistics usually involve some form of averaging, which increases the signal to noise (S/N) and can motivate assuming an approximately Gaussian likelihood when performing inference (e.g. Greig & Mesinger 2015; Gazagnes et al. 2021; Watkinson et al. 2022). [^3] The gaussianity of summaries constructed by averaging can be loosely motivated by appealing to the Central Limit Theorem. However, the theorem holds only for independent, identically distributed random variables. Unfortunately, we only have a single observable Universe; thus summaries of cosmological data sets (e.g. the power spectrum) often involve binning over correlated (i.e. not independent) and/or differently evolving fields (i.e. not identically distributed), resulting in non-Gaussian likelihoods (e.g. Prelogović & Mesinger 2023).
@@
The most common choice of a 21-cm summary statistic is the spherically-averaged power spectrum (1D PS). The 1D PS maximizes the signal-to-noise (S/N) of an interferometric observation – a primary concern in the run up to a first detection (e.g. Greig & Mesinger 2018; Trott et al. 2020; Mertens et al. 2020; HERA Collaboration et al. 2023). Another common choice of a summary is the cylindrically-averaged (2D) PS, which includes additional information on the anisotropy of the cosmic 21-cm signal (e.g. Bharadwaj & Ali 2005; Barkana & Loeb 2006; Datta et al. 2012; Mao et al. 2012). Fourier space is also a natural basis to isolate the cosmic signal from instrumental effects and foregrounds, which primarily reside in a wedge region in the 2D PS (e.g. Morales et al. 2012; Vedantham et al. 2012; Trott et al. 2012; Parsons et al. 2014; Liu et al. 2014a, b; Murray & Trott 2018). However, the power spectrum ignores Fourier phases, which can encode significant information for a highly non-Gaussian signal such as 21-cm from the EoR/CD (e.g. Bharadwaj & Pandey 2005; Mellema et al. 2006; Shimabukuro et al. 2016; Majumdar et al. 2018; Watkinson et al. 2019). Because of this, several studies have also explored non-Gaussian statistics, such as the bispectrum (Shimabukuro et al., 2016; Majumdar et al., 2018; Watkinson et al., 2019, 2022), trispectrum (Cooray et al., 2008; Flöss et al., 2022), morphological spectra (Gazagnes et al., 2021), bubble size distribution (Lin et al., 2016; Giri et al., 2018; Shimabukuro et al., 2022; Doussot & Semelin, 2022; Lu et al., 2023), one-point statistics (Harker et al., 2009; Watkinson & Pritchard, 2015; Kittiwisit et al., 2022), genus topology (Hong et al., 2014), Wavelet Scattering Transform (Greig et al., 2022, 2023; Zhao et al., 2023a), Minkowski functionals and tensors (Gleser et al., 2006; Yoshiura et al., 2017; Chen et al., 2019; Kapahtia et al., 2019; Spina et al., 2021) and Betti numbers (Giri & Mellema, 2021; Kapahtia et al., 2021).
@@
Is there an ?optimal? choice of summary statistic? One way to define ?optimal? is by how tightly we can recover cosmological/astrophysical parameters. This can be quantified through the Fisher information matrix (Fisher, 1935). The Fisher information is commonly used to make forecasts using pre-determined summary statistics, or even to define the summary statistic itself. Optimal summaries/algorithms that are based on the Fisher matrix include Karhunen-Loève methods and Massively Optimised Parameter Estimation and Data compression (MOPED) for linear transformations (e.g. Tegmark et al. 1997; Heavens et al. 2000), as well as non-linear generalizations like Information Maximizing Neural Networks (IMNNs, Charnock et al. 2018) and Fishnets (Makinen et al., 2023). However, these summary statistics come at the cost of interpretability, making the compression more of a ?black box? compared to physically-motivated and easy to interpret summaries like the PS. [^4] An interesting question is why would we even expect a truly optimal (lossless) compression to exist? The answer to this can be motivated through the manifold hypothesis (Fefferman et al., 2016), which states that physical probability distributions of the data often lie on a low-dimensional manifold. Optimal (lossless) compression would then ?extract? such manifolds from the original high-dimensional data. The above-mentioned algorithms do not attempt this directly, but aim at making compression as optimal as possible, given certain assumptions.
@@
It is also important to note that, unlike $\Lambda$CDM, there is no obvious or unique choice to parametrize the astrophysics of early galaxies whose radiation determines the cosmic 21-cm signal. Neither do we have a good idea of a ?fiducial? model. Thus there is no guarantee that a summary statistic that maximizes Fisher information at a single parameter value for a single model parametrization would generalize well to other parameters / models.
@@
Here we systematically compare the information content of several common 21-cm summaries on the basis of their Fisher information. These include: (i) the spherically-averaged power spectrum (1DPS), (ii) the cylindrically-averaged power spectrum (2DPS), (iii) the 2D Wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; and (v) an IMNN. Several previous studies compared the constraining power of 21-cm summaries (e.g. Gazagnes et al. 2021; Watkinson et al. 2022; Greig et al. 2022; Zhao et al. 2023b; Hothi et al. 2023). Here we explore a broader range of summary statistics, introducing also IMNNs to the field. Unlike previous 21-cm Fisher studies, we do not make the simplifying assumption of a constant covariance matrix. Most importantly, we compare the summary statistics across a broad range of parameter space, instead of choosing only one or two fiducial values to make a mock 21-cm observation. This is very important to verify the robustness of the summary, since we do not have a good idea of a ?fiducial? astrophysical model for the first galaxies.
@@
The paper is organized as follows. In section §2 we define the information content of a summary through the Fisher matrix and introduce the Fisher information distribution. In section §3 we introduce our 21-cm simulator and explain the summaries we consider in detail. In §4 we discuss our database of prior samples for evaluating the Fisher information. In §5 we present our results, quantifying the most informative summaries as well as the error introduced by the common assumption of a constant covariance matrix. Finally, we present concluding remarks and future prospects in §6. All quantities are quoted in co-moving units, and we assume a standard $\Lambda\mathrm{CDM}$ cosmology: $\left(\Omega_{\Lambda},\Omega_{\mathrm{M}},\Omega_{b},n,\sigma_{8},H_{0}\right)=\left(0.69,0.31,0.048,0.97,0.81,68\mathrm{\leavevmode\nobreak\ km}\mathrm{\leavevmode\nobreak\ s}^{-1}\mathrm{Mpc}^{-1}\right)$, consistent with the results from Planck Collaboration et al. (2020).
@@ 2 The Fisher matrix as a measure of information content
How do we measure the information content of a summary? If we have a theoretical model with a given parametrization, we can see how sensitive the summary is to changes in the astrophysical/cosmological parameters around some fiducial value. Specifically, we can calculate the Fisher matrix (e.g. Spall 2005):
@@
where $\bm{d}$ is the data summary (c.f. Fig. 1), $P(\bm{d}|\bm{\theta})$ is the likelihood, and $\mathbb{E}$ denotes the expectation value over the likelihood (i.e. an empirical average over many realizations) evaluated at a given point in parameter space $\bm{\theta}^{*}$. The Fisher information matrix provides the maximum constraining power, known as the Cramer-Rao bound (Rao, 1992; Cramér, 1999): $\operatorname{Var}(\hat{\bm{\theta}}_{m})\geq(\bm{F}^{-1})_{mm}$, where $\hat{\bm{\theta}}$ is an unbiased estimator of the parameters. In the multivariate case, one can equivalently write:
@@
Therefore, the determinant of the Fisher matrix measures the tightest parameter volume that can be constrained by the data. From now on, we will refer to $\det\bm{F}(\bm{\theta}^{*})$ as simply the Fisher information around $\bm{\theta}^{*}$ the higher its value, the more constraining the summary is around $\bm{\theta}^{*}$.
@@ 2.1 Gaussian approximation
In almost all cosmological scenarios, we do not explicitly know the full likelihood function $P(\bm{d}|\bm{\theta})$, nor its ?score? $\nabla_{\bm{\theta}}\ln P(\bm{d}|\bm{\theta})$. Instead for performing quick forecasts, it is common to approximate the likelihood as a Gaussian in data space $P(\bm{d}|\bm{\theta})\approx\mathcal{N}(\bm{d}|\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$, where the mean $\bm{\mu}(\bm{\theta})$ and covariance $\bm{\Sigma}(\bm{\theta})$ are general functions of the parameter space.
@@
Here the first (second) term measures how the mean (covariance) of the data summary changes with respect to the parameters. Cosmological Fisher forecasts usually ignore one of the two terms (most often the second term $\bm{F}_{\bm{\Sigma}}$), effectively fixing either the mean or the covariance matrix to some fiducial value. However, we will see below that both terms are non-negligible for some summaries (e.g. Carron 2013).

### v2

# How informative are summaries of the cosmic 21 cm signal?

@@
How informative are summaries of the cosmic 21 cm signal?
@@
The cosmic 21 cm signal will bring data-driven advances to studies of Cosmic Dawn (CD) and the Epoch of Reionization (EoR). Radio telescopes such as the Square Kilometre Array (SKA) will eventually map the HI fluctuations over the first billion years – the majority of our observable Universe. With such large data volumes, it becomes increasingly important to develop ?optimal? summary statistics, which will allow us to learn as much as possible about the CD and EoR. In this work we compare the astrophysical parameter constraining power of several 21 cm summary statistics, using the determinant of the Fisher information matrix, $\det\bm{F}$. Since we do not have an established ?fiducial? model for the astrophysics of the first galaxies, we computed for each summary the distribution of $\det\bm{F}$ across the prior volume. Using a large database of cosmic 21 cm light cones that include realizations of telescope noise, we compared the following summaries: (i) the spherically averaged power spectrum (1DPS), (ii) the cylindrically averaged power spectrum (2DPS), (iii) the 2D wavelet scattering transform (WST), (iv) a recurrent neural network (RNN) trained as a regressor; (v) an information-maximizing neural network (IMNN); and (vi) the combination of 2DPS and IMNN. Our best performing individual summary is the 2DPS, which provides relatively high Fisher information throughout the parameter space. Although capable of achieving the highest Fisher information for some parameter choices, the IMNN does not generalize well, resulting in a broad distribution across the prior volume. Our best results are achieved with the concatenation of the 2DPS and IMNN. The combination of only these two complimentary summaries reduces the recovered parameter variances on average by factors of $\sim$6.5 – 9.5, compared with using each summary independently. Finally, we point out that that the common assumption of a constant covariance matrix when doing Fisher forecasts using 21 cm summaries can significantly underestimate parameter constraints.
@@ 1 Introduction
The cosmic 21 cm signal, corresponding to the spin flip transition of the ground state of HI, provides a window on the first billion years of the Universe’s evolution. This under-explored period witnessed fundamental cosmic milestones, including the Cosmic Dawn (CD) of the first galaxies and the final phase change of our Universe: the Epoch of Reionization (EoR). Current radio interferometers are setting increasingly tight upper limits on its power spectrum (PS; e.g., Trott et al. 2020; Mertens et al. 2020; Gehlot et al. 2020; HERA Collaboration et al. 2023; Munshi et al. 2024), while the upcoming Square Kilometre Array (SKA) [^1] https://www.skatelescope.org promises to provide a 3D image of the EoR in the next decade(s) (e.g., Koopmans et al. 2015; Mesinger 2020). The unknown properties of the first galaxies and intergalactic medium (IGM) structures are encoded in the timing and morphology of the cosmic 21 cm signal (e.g., Pritchard & Furlanetto 2007; McQuinn & O’Leary 2012; Visbal et al. 2012; Pacucci et al. 2014).
@@
The only statistically robust way to infer these properties is through Bayesian inference: sampling from theory and comparing forward models to observations. But what should we use to compare the forward model to the data?; in other words, how should we construct the likelidhood? The cosmic 21 cm signal is intrinsically a 3D light cone map. Performing inference directly on the light cone would be incredibly challenging due to the high dimensionality of the data. Effectively, at each frequency the SKA should obtain a sky map comparable to current cosmic microwave background maps, but with the full light cone including thousands of such frequency slices (e.g., Loeb & Zaldarriaga 2004). [^2] Inference on high-dimensional light cones is still possible in some cases, for example galaxy large-scale structure maps (e.g., Kitaura & Enßlin 2008; Jasche & Kitaura 2010; Jasche & Wandelt 2012, 2013; Leclercq et al. 2017; McAlpine et al. 2022; Dai & Seljak 2022; Bayer et al. 2023). Although such studies still compress the full galaxy observations, only treating them as biased tracers of the matter field, they are able to recover the phase information of our Universe’s initial conditions (so-called constrained realizations). It remains to be seen if such studies can be extended to the EoR and CD (though see Zhao et al. 2023b for a recent proof-of-concept study). Instead, the data are compressed into a summary statistic. Summary statistics usually involve some form of averaging, which increases the signal-to-noise ratio (S/N) and can motivate assuming an approximately Gaussian likelihood when performing inference (e.g., Greig & Mesinger 2015; Gazagnes et al. 2021; Watkinson et al. 2022). [^3] The Gaussianity of summaries constructed by averaging can be loosely motivated by appealing to the Central Limit Theorem. However, the theorem holds only for independent, identically distributed random variables. Unfortunately, we only have a single observable Universe; thus summaries of cosmological datasets (e.g., the PS) often involve binning over correlated (i.e., not independent) and/or differently evolving fields (i.e., not identically distributed), resulting in non-Gaussian likelihoods (e.g., Prelogović & Mesinger 2023).
@@
The most common choice of a 21 cm summary statistic is the spherically averaged power spectrum (1DPS). The 1DPS maximizes the S/N of an interferometric observation – a primary concern in the run-up to a first detection (e.g., Greig & Mesinger 2018; Trott et al. 2020; Mertens et al. 2020; HERA Collaboration et al. 2023). Another common summary choice is the cylindrically averaged power spectrum (2DPS), which includes additional information on the anisotropy of the cosmic 21 cm signal (e.g., Bharadwaj & Ali 2005; Barkana & Loeb 2006; Datta et al. 2012; Mao et al. 2012). Fourier space is also a natural basis for isolating the cosmic signal from instrumental effects and foregrounds, which primarily reside in a wedge region in the 2DPS (e.g., Morales et al. 2012; Vedantham et al. 2012; Trott et al. 2012; Parsons et al. 2014; Liu et al. 2014a, b; Murray & Trott 2018). However, the PS ignores Fourier phases, which can encode significant information for a highly non-Gaussian signal such as the 21 cm signal from the EoR and CD (e.g., Bharadwaj & Pandey 2005; Mellema et al. 2006; Shimabukuro et al. 2016; Majumdar et al. 2018; Watkinson et al. 2019). Because of this, several studies have also explored non-Gaussian statistics, such as the bispectrum (Shimabukuro et al., 2016; Majumdar et al., 2018; Watkinson et al., 2019, 2022), the trispectrum (Cooray et al., 2008; Flöss et al., 2022), morphological spectra (Gazagnes et al., 2021), the bubble size distribution (Lin et al., 2016; Giri et al., 2018; Shimabukuro et al., 2022; Doussot & Semelin, 2022; Lu et al., 2024), one-point statistics (Harker et al., 2009; Watkinson & Pritchard, 2015; Kittiwisit et al., 2022), genus topology (Hong et al., 2014), wavelet scattering transform (WST; Greig et al., 2022, 2023; Zhao et al., 2023a), Minkowski functionals and tensors (Gleser et al., 2006; Yoshiura et al., 2017; Chen et al., 2019; Kapahtia et al., 2019; Spina et al., 2021), and Betti numbers (Giri & Mellema, 2021; Kapahtia et al., 2021).
@@
Is there an ?optimal? choice of summary statistic? One way to define optimal is by how tightly we can recover cosmological and astrophysical parameters. This can be quantified through the Fisher information matrix (Fisher, 1935). The Fisher information is commonly used to make forecasts using predetermined summary statistics, or even to define the summary statistic itself. Optimal summaries (algorithms) that are based on the Fisher matrix include Karhunen-Loève methods and massively optimized parameter estimation and data compression for linear transformations (e.g., Tegmark et al. 1997; Heavens et al. 2000), as well as nonlinear generalizations like information-maximizing neural networks (IMNNs; Charnock et al. 2018) and fishnets (Makinen et al., 2023). However, these summary statistics come at the cost of interpretability, making the compression more of a ?black box? compared to physically motivated and easy-to-interpret summaries like the PS. [^4] An interesting question is why would we even expect a truly optimal (lossless) compression to exist? The answer to this can be motivated through the manifold hypothesis (Fefferman et al., 2016), which states that physical probability distributions of the data often lie on a low-dimensional manifold. Optimal (lossless) compression would then ?extract? such manifolds from the original high-dimensional data. The above-mentioned algorithms do not attempt this directly, but aim at making compression as optimal as possible, given certain assumptions.
@@
It is also important to note that, unlike $\Lambda$ cold dark matter ($\Lambda$CDM), there is no obvious or unique choice for parameterizing the astrophysics of early galaxies whose radiation determines the cosmic 21 cm signal. Nor do we have a good idea of a ?fiducial? model. Thus, there is no guarantee that a summary statistic that maximizes Fisher information at a single parameter value for a single model parametrization would generalize well to other parameters and models.
@@
Here we systematically compare the information content of several common 21 cm summaries on the basis of their Fisher information. These include: (i) the 1DPS, (ii) the 2DPS, (iii) the 2D WST, (iv) a recurrent neural network (RNN) trained as a regressor; and (v) an IMNN.555 Note that the dimensionality of the PS (1D or 2D) labels the dimensionality of the summary, while for the WST, 2D labels the dimensionality of the data on which it is computed. In our case, those are 2D sky-plane images. Several previous studies compared the constraining power of 21 cm summaries (e.g., Gazagnes et al. 2021; Watkinson et al. 2022; Greig et al. 2022; Zhao et al. 2023b; Hothi et al. 2023). Here we explore a broader range of summary statistics, introducing IMNNs to the field. Unlike previous 21 cm Fisher studies, we do not make the simplifying assumption of a constant covariance matrix. Most importantly, we compare the summary statistics across a broad parameter space, instead of choosing only one or two fiducial values to make a mock 21 cm observation. This is very important for verifying the robustness of the summary since we do not have a fiducial astrophysical model for the first galaxies.
@@
The paper is organized as follows. In Sect. 2 we define the information content of a summary through the Fisher matrix and introduce the Fisher information distribution. In Sect. 3 we introduce our 21 cm simulator and explain the summaries we consider in detail. In Sect. 4 we discuss our database of prior samples for evaluating the Fisher information. In Sect. 5 we present our results, quantifying the most informative summaries as well as the error introduced by the common assumption of a constant covariance matrix. Finally, we present concluding remarks and future prospects in Sect. 6. All quantities are quoted in comoving units, and we assume a standard $\Lambda\mathrm{CDM}$ cosmology: $\left(\Omega_{\Lambda},\Omega_{\mathrm{M}},\Omega_{b},n,\sigma_{8},H_{0}\right)=\left(0.69,0.31,0.048,0.97,0.81,68\mathrm{\leavevmode\nobreak\ km}\mathrm{\leavevmode\nobreak\ s}^{-1}\mathrm{Mpc}^{-1}\right)$, consistent with the results from Planck Collaboration et al. (2020).
@@ 2 The Fisher matrix as a measure of information content
How do we measure the information content of a summary? If we have a theoretical model with a given parametrization, we can see how sensitive the summary is to changes in the astrophysical/cosmological parameters around some fiducial value. Specifically, we can calculate the Fisher matrix (e.g., Spall 2005):
@@
where $\bm{d}$ is the data summary (c.f. Fig. 1), $P(\bm{d}|\bm{\theta})$ is the likelihood, and $\mathbb{E}$ denotes the expectation value over the likelihood (i.e., an empirical average over many realizations) evaluated at a given point in parameter space $\bm{\theta}^{*}$. The Fisher information matrix provides the maximum constraining power, known as the Cramér-Rao bound (Rao, 1992; Cramér, 1999): $\operatorname{Var}(\hat{\bm{\theta}}_{m})\geq(\bm{F}^{-1})_{mm}$, where $\hat{\bm{\theta}}$ is an unbiased estimator of the parameters. In the multivariate case, one can equivalently write:
@@
Therefore, the determinant of the Fisher matrix measures the tightest parameter volume that can be constrained by the data. From now on, we refer to $\det\bm{F}(\bm{\theta}^{*})$ as simply the Fisher information around $\bm{\theta}^{*}$ — the higher its value, the more constraining the summary is around $\bm{\theta}^{*}$.
@@ 2.1 Gaussian approximation
In almost all cosmological scenarios, we do not explicitly know the full likelihood function $P(\bm{d}|\bm{\theta})$, nor its ?score,? $\nabla_{\bm{\theta}}\ln P(\bm{d}|\bm{\theta})$. For performing quick forecasts, it is instead common to approximate the likelihood as a Gaussian in data space $P(\bm{d}|\bm{\theta})\approx\mathcal{N}(\bm{d}|\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$, where the mean $\bm{\mu}(\bm{\theta})$ and covariance $\bm{\Sigma}(\bm{\theta})$ are general functions of the parameter space.
@@
Here the first (second) term measures how the mean (covariance) of the data summary changes with respect to the parameters. Cosmological Fisher forecasts usually ignore one of the two terms (most often the second term $\bm{F}_{\bm{\Sigma}}$), effectively fixing either the mean or the covariance matrix to some fiducial value. However, we will see below that both terms are non-negligible for some summaries (e.g., Carron 2013).